A card chosen at random from a standard deck is not an 8 or an ace.
Answer by dyakobovitch(40)

Our objective is to add up these probabilities and subtract from 1 to find the probability that we are neither an 8 or an ace.

The probability of getting an ace is 4/52=1/13. The same goes for an 8. So, our probability is 1/13 + 1/13 = 2/13.

The probability of neither an 8 nor an ace is 1-2/13 = 11/13.

Feel free to leave any comments or concerns. -- David